# LocoTouch: Learning Dexterous Quadrupedal Transport with Tactile Sensing

### v1

@@ D.2 Reward Design
Object xy position $\|\bm{p}_{\text{o},xy}^{\text{w}}-\bm{p}_{xy}^{\text{w}}\|_{2}$ $-50.0$
Object yaw angle $\|\bm{\theta}_{\text{o},z}^{\text{r}}\|_{2}$ $-0.1$
Foot dragging $\sum_{i=1}^{4}\mathbf{1}_{\left\{\bm{p}_{\text{f}_{i},z}^{\text{w}}\leq h_{z,\text{th}}\land\|\bm{v}_{\text{f}_{i},xy}^{\text{w}}\|^{2}\geq v_{xy,\text{th}}\right\}}$ $-1.0$
Foot slipping $\sum_{i=1}^{4}\|\bm{v}_{\text{f}_{i},xy}^{\text{w}}\|\cdot\mathbf{1}_{\{\bm{F}_{\text{f}_{i},z}^{\text{w}}\geq F_{z,\text{th}}\}}$ $-0.1$
@@
Object z velocity $\|\bm{v}_{\text{o},z}^{\text{r}}\|_{2}$ $-0.5$
Object roll angle and velocity $\|\bm{\theta}_{\text{o},x}^{\text{r}}\|_{2}+\|\bm{w}_{\text{o},x}^{\text{r}}\|_{2}$ $-0.05$
Object dangerous state $\mathbf{1}_{\{\text{danger}\}}$ $-50.0$
Base height tracking $|\bm{p}_{z}^{\text{w}}-h_{\text{target}}|^{2}$ $-0.5$
Base z velocity $|\bm{v}_{z}^{\text{w}}|^{2}$ $-1.0$
Base roll-pitch angle $|\bm{\theta}_{x}^{\text{w}}|^{2}+|\bm{\theta}_{y}^{\text{w}}|^{2}$ $-1.0$
Base roll-pitch velocity $|\bm{w}_{x}^{\text{r}}|+|\bm{w}_{y}^{\text{r}}|$ $-0.2$
Joint position deviation $\|\bm{q}-\bm{q}_{\text{default}}\|_{2}\cdot\left(\mathbf{1}_{\{\text{move}\}}+\alpha_{\text{stance}}\,\mathbf{1}_{\{\text{stand}\}}\right)$ $-0.5$
Joint position limits $\sum_{j=1}^{12}[q^{\min}_{j}-q_{j}]_{+}+[q_{j}-q^{\max}_{j}]_{+}$ $-10.0$
Joint velocity $\|\dot{\bm{q}}\|_{2}$ $-5\mathrm{e}{-3}$
Joint acceleration $\|\ddot{\bm{q}}\|_{2}$ $-5\mathrm{e}{-6}$
Joint torque $\|\bm{\tau}\|_{2}$ $-2.5\mathrm{e}{-4}$
Action rate $\sum_{j=1}^{12}\|\mathbf{q}_{j,\text{target}}-\mathbf{q}_{j,\text{last target}}\|$ $-0.75$
Thigh–calf collision $\sum_{i=1}^{4}\mathbf{1}_{\{\|\bm{F}_{t_{i}}\|_{2}>f_{\text{th}}\}}+\mathbf{1}_{\{\|\bm{F}_{c_{i}}\|_{2}>f_{\text{th}}\}}$ $-5.0$
@@
Learning epochs per training batch $5$
Number of mini-batches per training batch $4$
where $f_{i}$ is the $i$-th foot of the robot, $p^{w}_{i,z}$ is its height above the ground, $\mathbf{v}^{w}_{i,xy}$ is the lateral linear velocity of foot $i$ in the world frame, $h_{z,\text{th}}$ is the height threshold, and $v_{xy,\text{th}}$ is the lateral velocity threshold. In our implementation, we set $h_{z,\text{th}}$ to $0.015\leavevmode\nobreak\ \mathrm{m}$ and $v_{xy,\text{th}}$ to $0.1\leavevmode\nobreak\ \mathrm{m/s}$.
D.3 Symmetricity Function
@@
D.4 Details for Tactile Policy Distillation
We summarize the tactile policy structure in Table 8 and the hyper-parameters for distillation in Table 8.
@@ D.4 Details for Tactile Policy Distillation
Appendix E Experiment Details
E.1 Diverse Objects Transport
@@
E.2 Long-Distance Transport
We demonstrate that LocoTouch is capable of transporting objects over long distances, even under varying velocity commands and sharp turns. As shown in Figure 17, after carrying the slippery drink bottle for approximately $60\mathrm{m}$, the robot continues to maintain the object near the center of its back. This highlights the robustness of LocoTouch during real-world deployment. Videos are available on the website page.
@@ E.2 Long-Distance Transport
E.3 Performance of the Locomotion-Only Policy

### v2

@@ C.4 Reward Design
Object xy position $\|\bm{p}_{\text{o},xy}^{\text{w}}-\bm{p}_{xy}^{\text{w}}\|_{2}$ 50.0-50.0
Object yaw angle $\|\bm{\theta}_{\text{o},z}^{\text{r}}\|_{2}$ 0.1-0.1
Foot dragging $\sum_{i=1}^{4}\mathbf{1}_{\left\{\bm{p}_{\text{f}_{i},z}^{\text{w}}\leq h_{z,\text{th}}\land\|\bm{v}_{\text{f}_{i},xy}^{\text{w}}\|^{2}\geq v_{xy,\text{th}}\right\}}$ 1.0-1.0
Foot slipping $\sum_{i=1}^{4}\|\bm{v}_{\text{f}_{i},xy}^{\text{w}}\|\cdot\mathbf{1}_{\{\bm{F}_{\text{f}_{i},z}^{\text{w}}\geq F_{z,\text{th}}\}}$ 0.1-0.1
@@
Object z velocity $\|\bm{v}_{\text{o},z}^{\text{r}}\|_{2}$ 0.5-0.5
Object roll angle and velocity $\|\bm{\theta}_{\text{o},x}^{\text{r}}\|_{2}+\|\bm{w}_{\text{o},x}^{\text{r}}\|_{2}$ 0.05-0.05
Object dangerous state $\mathbf{1}_{\{\text{danger}\}}$ 50.0-50.0
Base height tracking $|\bm{p}_{z}^{\text{w}}-h_{\text{target}}|^{2}$ 0.5-0.5
Base z velocity $|\bm{v}_{z}^{\text{w}}|^{2}$ 1.0-1.0
Base roll-pitch angle $|\bm{\theta}_{x}^{\text{w}}|^{2}+|\bm{\theta}_{y}^{\text{w}}|^{2}$ 1.0-1.0
Base roll-pitch velocity $|\bm{w}_{x}^{\text{r}}|+|\bm{w}_{y}^{\text{r}}|$ 0.2-0.2
Joint position deviation $\|\bm{q}-\bm{q}_{\text{default}}\|_{2}\cdot\left(\mathbf{1}_{\{\text{move}\}}+\alpha_{\text{stance}}\,\mathbf{1}_{\{\text{stand}\}}\right)$ 0.5-0.5
Joint position limits $\sum_{j=1}^{12}[q^{\min}_{j}-q_{j}]_{+}+[q_{j}-q^{\max}_{j}]_{+}$ 10.0-10.0
Joint velocity $\|\dot{\bm{q}}\|_{2}$ $5\mathrm{e}{-3}$
Joint acceleration $\|\ddot{\bm{q}}\|_{2}$ $5\mathrm{e}{-6}$
Joint torque $\|\bm{\tau}\|_{2}$ $2.5\mathrm{e}{-4}$
Action rate $\sum_{j=1}^{12}\|\mathbf{q}_{j,\text{target}}-\mathbf{q}_{j,\text{last target}}\|$ 0.75-0.75
Thigh–calf collision $\sum_{i=1}^{4}\mathbf{1}_{\{\|\bm{F}_{t_{i}}\|_{2}>f_{\text{th}}\}}+\mathbf{1}_{\{\|\bm{F}_{c_{i}}\|_{2}>f_{\text{th}}\}}$ 5.0-5.0
@@
Learning epochs per training batch 55
Number of mini-batches per training batch 44
C.5 Symmetricity Function
@@
Appendix D Details for Tactile Policy Distillation
During distillation, we disable the velocity curriculum and set the zero-command step number $k=50$. We summarize the tactile policy structure in Table 10 and the hyper-parameters for distillation in Table 10.
@@ Appendix D Details for Tactile Policy Distillation
Appendix E Supplementary Experiments in the Real World
E.1 System Setup
Figure 15: System setup in the real world.
E.2 Diverse Cylindrical Objects Transport
@@
E.3 Generalization to Non-Cylindrical Everyday Objects
Our experiments primarily focus on cylindrical objects, as they represent the most challenging case for transport. Nevertheless, we also evaluate LocoTouch on a variety of non-cylindrical everyday objects, such as cups and wrenches (Figure 18), and find that it can stably balance and transport them. This indicates that LocoTouch does not overfit to the contact patterns of cylindrical objects, likely benefiting from the diverse and unstructured contact patterns generated by the dynamic interactions from cylinders during training. Notably, when we place a cup slightly off-center on the robot’s back, LocoTouch is able to actively adjust it toward the center during transport. Corresponding videos are available on the project website.
Figure 18: LocoTouch balances and transports everyday objects such as cups and wrenches.
E.4 Robustness in Uneven Terrains
Although LocoTouch is trained only on flat terrain in simulation, we find that it generalizes well to diverse environments, including slopes, gravel, and rough surfaces (Fig. 19). This robustness is likely attributed to the simulated perturbations and domain randomization introduced during training. Nevertheless, we believe that incorporating visual perception represents a promising direction to further enhance reliability in more challenging terrains.
Figure 19: LocoTouch generalizes to diverse uneven terrains, including slopes, gravel, and rough surfaces.
E.5 Long-Distance Transport
We demonstrate that LocoTouch is capable of transporting objects over long distances, even under varying velocity commands and sharp turns. As shown in Figure 20, after carrying the slippery drink bottle for approximately $60\mathrm{m}$, the robot continues to maintain the object near the center of its back. This highlights the robustness of LocoTouch during real-world deployment.
@@ E.5 Long-Distance Transport
E.6 Performance of the Locomotion-Only Policy